OpenStudy (anonymous):

The ratio of red marbles to blue marbles in a jar is 3 to 8. There are 40 blue marbles in the jar. How many red marbles are in the jar? A. 5 B. 15 C. 24 D. 35

OpenStudy (anonymous):

@4jesus

OpenStudy (anonymous):

@ccswims

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

Do I multiply 3x8=24... Then 40-24

OpenStudy (anonymous):

Nvm that'll equal 16!

OpenStudy (anonymous):

Good job!

OpenStudy (kropot72):

First find out how many lots of 8 blue marbles are in the 40 blue marbles by dividing 40 by 8. Then multiply 3 by the result: \[\large Number\ of \ red\ marbles=\frac{40}{8}\times3=you\ can\ calculate\]

OpenStudy (michele_laino):

please note that you have to use this formula: \[\frac{ r }{ b }=\frac{ 3 }{ 8 }\] where r is the number of red Marbles, and b is the number of blue Marbles, furthermore, we know that: \[r+b=40\]

OpenStudy (anonymous):

B

OpenStudy (anonymous):

40/8=5. Then 5*3=15 :)

OpenStudy (michele_laino):

oops.. furthermor we have b= 40

OpenStudy (anonymous):

Thanks guys

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

@Michele_Laino People makes mistakes. I'm happy you helped though. :)

OpenStudy (michele_laino):

thanks! I was too hurry!